### v2

@@ Abstract
The recent work of Horikawa and Nakagawa, (2024) claims that under a complete market admitting statistical arbitrage, the difference between the hedging position provided by deep hedging and that of the replicating portfolio is a statistical arbitrage. This raises concerns as it entails that deep hedging can include a speculative component aimed simply at exploiting the structure of the risk measure guiding the hedging optimisation problem. We test whether such finding remains true in a GARCH-based market model. We observe that the difference between deep hedging and delta hedging can be a statistical arbitrage if the risk measure considered does not put sufficient relative weight on adverse outcomes. Nevertheless, a suitable choice of risk measure can prevent the deep hedging agent from including a speculative overlay within its hedging strategy.
@@ 1   Introduction
The seminal paper of Buehler et al., (2019), which proposes to use deep reinforcement learning (RL) methods to obtain optimal hedging procedures for financial derivatives, initiated a whole strand of literature: Halperin, (2019), Cao et al., (2020), Du et al., (2020), Carbonneau and Godin, (2021), Carbonneau, (2021), Horvath et al., (2021), Imaki et al., (2021), Lütkebohmert et al., (2022), Cao et al., (2023), Carbonneau and Godin, (2023), Marzban et al., (2023), Mikkilä and Kanniainen, (2023), Raj et al., (2023) and Wu and Jaimungal, (2023). See also Hambly et al., (2023) and Pickard and Lawryshyn, (2023) for related surveys. Deep RL methods are particulary well-suited to solve dynamic hedging problems as such methods can handle the curse of dimensionality, a problem that more traditional methods (e.g., finite elements dynamic programming) struggle to overcome. They can also work with very general dynamics for hedging asset prices, not being limited by mathematical tractability issues.
@@
While the ability of deep hedging strategies to outperform standard counterparts is well-documented, the existing literature has not yet extensively analyzed the structure of optimal policies and explained how such incremental performance is attained. Neagu et al., (2024) make a step in that direction by investigating the impact of the various features on the optimal hedging decision in the presence of illiquidity market impacts.
@@
In their recent work, Horikawa and Nakagawa, (2024) investigate complete markets that allow for statistical arbitrage with respect to a specific risk measure $\rho$. They assert that, within this framework, hedging strategies that minimize the chosen risk metric combine the traditional delta-hedging approach with a statistical arbitrage overlay. In a more general market setting, Buehler et al., (2021) find that the optimal hedging strategy maximizing the entropy utility can also incorporate a statistical arbitrage component. Such claims raise concern about the suitability of the deep hedging approach; incorporating speculative or arbitrage-like components that do not contribute to reducing the risk exposure within hedging portfolios would be deemed undesirable in practice. Our objective is therefore to assess empirically whether deep hedging policies minimizing conventional risk metrics contain a speculative component in a GARCH-based market setting.
@@
The paper is divided as follows. Section 2 provides the hedging problem formulation. Section 3 describes the deep hedging framework used to solve the problem, and discusses the common delta hedging benchmark. Numerical experiments assessing the behavior of the deep versus delta hedging difference strategy are provided in Section 4.111The Python code which allows replicating the numerical experiments from this paper can be found at https://github.com/cpmendoza/DeepHedging_StatisticalArbitrage.git. Section 5 concludes.
@@ 2   Market model for hedging
This paper considers dynamic hedging strategies for European call options, which involve the construction of a self-financing hedging portfolio composed of the underlying asset $S$ and a cash account. The portfolio is rebalanced daily to optimally offset the net risk exposure at the option maturity, denoted as $T$ days. The time-$t$ underlying asset price is $S_{t}$. The hedging strategy is represented by the predictable process $\delta=\{\delta_{t}\}_{t=1}^{T}$, where $\delta_{t}$ is the number of underlying asset shares held during the interval $(t-1,t]$. The time-$t$ discounted gain made by the hedging portfolio is $G_{t}^{\delta}=\sum_{k=1}^{t}\delta_{k}(\beta^{k}S_{k}\mbox{e}^{q\Lambda}-\beta^{k-1}S_{k-1})$ with $\beta=\mbox{e}^{-r\Lambda}$, where $r$ is the annualized continuously compounded risk-free rate, $q$ is the annualized underlying asset dividend yield, and the period length is $\Lambda=\frac{1}{252}$ years. The time-$t$ self-financing hedging portfolio value is
@@
where $\rho$ is the risk measure used by the hedging agent to quantify risk. In this paper we consider the Conditional Value-at-Risk (CVaRα) defined as $\rho(\xi_{T}^{\delta})=\mathbb{E}[\xi_{T}^{\delta}\mid\xi_{T}^{\delta}\geq\mbox{VaR}_{\alpha}(\xi_{T}^{\delta})]$, where $\alpha\in(0,1)$ and $\mbox{VaR}_{\alpha}(\xi_{T}^{\delta})$ is the Value-at-Risk defined as $\mbox{VaR}_{\alpha}(\xi_{T}^{\delta})=\min_{c}\{c:\mathbb{P}(\xi_{T}^{\delta}\leq c)\geq\alpha\}$.
@@ 3.1   Deep hedging
The deep hedging (DH) framework, introduced by Buehler et al., (2019), provides a solution to the hedging problem (3) by leveraging reinforcement learning (RL) techniques. The DH policy $\tilde{\delta}$ is approximated with a neural network $\delta^{DH}_{\theta}$ with parameters $\theta$, which returns a hedging position $\delta_{t+1}$ when provided with time-$t$ input features $X_{t}$. The objective function to be minimized is thus
@@
which ensures that the cash amount borrowed in the hedging portfolio remains below $B>0$ (see François et al., (2024)).
@@ 3.3   Statistical arbitrage
which entails a trading strategy with zero initial investment that is deemed strictly less risky than a null investment according to risk measure $\rho$. Such definition is also in line with that of Buehler et al., (2021), who focus on the case of the entropy risk measure.
@@
is a statistical arbitrage strategy according to risk measure $\rho$. We wish to further extend their study and examine if the trading strategy $\delta^{-}$ behaves like a statistical arbitrage in a GARCH-based market model. In other words, we investigate whether the deep hedging approach typically incorporates a speculative arbitrage-like component aimed at exploiting the structure of the risk measure considered.
@@ 4.1   Stochastic market dynamics
where $\mu,\,\gamma\in\mathbb{R}$, $\omega,\,\alpha,\,\beta$ are positive, $\mathds{1}_{A}$ is the binary dummy variable indicating if event $A$ occurs and $\{\epsilon_{t}\}_{t=1}^{T}$ are independent standard normal random variables. Parameter estimates are obtained through maximum likelihood on a daily time series of the S&P 500 index extending from January 4, 2016, to December 31, 2020. Estimated parameters are $\mu=0.06\%$, $\omega=0.01\%$, $\alpha=0.11$, $\gamma=0.20$ and $\beta=0.78$. Furthermore, in all experiments, the annualized continuously compounded risk-free rate and dividend yield are assumed to be constant with values set to $r=2.66\%$ and $q=1.77\%$, respectively.
@@ 4.2   Comparative analysis of deep hedging and delta hedging strategies
In this section, we study the relationship between delta hedging and deep hedging. More specifically, we investigate whether the difference between both strategies represents a speculative overlay reminiscent of a statistical arbitrage. The comparison is conducted by hedging at-the-money (ATM) European call options, with $S_{0}=K=100$. The option maturity is $T=63$ days. The leverage constraint is $B=100$.
@@
For all confidence levels $\alpha$ below $50\%$, the strategy $\delta^{-}$ exhibits both positive average profitability $\mathbb{E}[V_{T}^{\delta^{-}}(0)]$ and a CVaR value $\rho(-V_{T}^{\delta^{-}}(0))$ that is negative. This corresponds to a formal statistical arbitrage strategy. Moreover, for the case $\alpha=50\%$, even if the risk measure is positive, it is nevertheless negligible in comparison to expected profits. The strategy $\delta^{-}$, though not a statistical arbitrage from the definition, exhibits a behavior that is quite similar to a statistical arbitrage. Conversely, difference strategies $\delta^{-}$ using $\alpha\geq 85\%$ clearly do not qualify as statistical arbitrage; the associated risk measure is high and the average profitability is negative.
@@
  • Results are for a short position on an ATM call option with a maturity of $N=63$ days. They are computed using 100,000 out-of-sample paths. The metric $\varrho$ denotes the (unconditional) Spearman correlation between underlying asset positions of the delta hedging strategy and the deep hedging strategy across all rebalancing days while $R^{2}$ represents the $R^{2}$ statistic obtained after regressing deep hedging positions onto delta hedging positions.
@@ 5   Conclusion
The two main conclusions from this study are therefore that (i) the objective function of the deep hedging problem must be carefully selected to prevent the hedging agent from abandoning its hedging objective and pursuing speculative behavior, and (ii) deep hedging can soundly achieve its hedging objectives when provided with a suitable risk measure.

### v3

@@ Abstract
The recent work of Horikawa and Nakagawa, (2024) claims that under a complete market admitting statistical arbitrage, the difference between the hedging position provided by deep hedging and that of the replicating portfolio is a statistical arbitrage. This raises concerns as it entails that deep hedging can include a speculative component aimed simply at exploiting the structure of the risk measure guiding the hedging optimisation problem. We test whether such finding remains true in a GARCH-based market model, which is an illustrative case departing from complete market dynamics. We observe that the difference between deep hedging and delta hedging is a speculative overlay if the risk measure considered does not put sufficient relative weight on adverse outcomes. Nevertheless, a suitable choice of risk measure can prevent the deep hedging agent from engaging in speculation.
@@ 1   Introduction
The seminal paper of Buehler et al., (2019), which proposes to use deep reinforcement learning (RL) methods to obtain optimal hedging procedures for financial derivatives, initiated a recent strand of literature. [^1] See for instance Halperin, (2019), Cao et al., (2020), Du et al., (2020), Carbonneau and Godin, (2021), Carbonneau, (2021), Horvath et al., (2021), Imaki et al., (2021), Lütkebohmert et al., (2022), Cao et al., (2023), Carbonneau and Godin, (2023), Marzban et al., (2023), Mikkilä and Kanniainen, (2023), Raj et al., (2023) and Wu and Jaimungal, (2023). See also Hambly et al., (2023) and Pickard and Lawryshyn, (2023) for related surveys. Deep RL methods are particulary well-suited to solve dynamic hedging problems because these methods can handle the curse of dimensionality, a problem that more traditional approaches (e.g., finite elements dynamic programming) struggle to overcome. They can also work with very general dynamics for asset prices, not being limited by mathematical tractability issues.
@@
While the ability of deep hedging strategies to outperform standard counterparts is well-documented, the existing literature has not yet extensively analyzed the structure of optimal policies and explained how such incremental performance is attained. Neagu et al., (2024) make a step in that direction by investigating the impact of the various features on optimal risk management decisions in the presence of illiquidity market impacts.
@@
In their recent work, Horikawa and Nakagawa, (2024) investigate complete markets that allow for statistical arbitrage with respect to a specific risk measure $\rho$. They assert that, within this framework, deep hedging strategies that minimize the chosen risk metric combine the traditional delta-hedging approach with a statistical arbitrage overlay. In a vector auto-regressive stochastic volatility model and in a GAN-simulated market model, Buehler et al., (2021) find that the optimal hedging strategy maximizing the entropy utility can also incorporate a statistical arbitrage component. Such claims raise concerns about the suitability of the deep hedging approach; incorporating speculative or arbitrage-like components that do not contribute to reducing the risk exposure within hedging portfolios would be deemed undesirable in practice. Our objective is therefore to assess empirically whether deep hedging policies minimizing conventional risk metrics still contain a speculative component in incomplete market settings, which would generalize the complete market conclusion of Horikawa and Nakagawa, (2024). We use a GARCH-based market setting as an illustrative example.
@@
The paper is divided as follows. Section 2 provides the hedging problem formulation. Section 3 describes the deep hedging framework used to solve the problem, and discusses the delta hedging benchmark. Numerical experiments assessing the behavior of the deep versus delta hedging difference strategy are provided in Section 4.222The Python code which allows replicating the numerical experiments from this paper can be found at https://github.com/cpmendoza/DeepHedging_StatisticalArbitrage.git. Section 5 concludes.
@@ 2   Market model for hedging
This paper considers dynamic risk management strategies for European call options, which involve the construction of a self-financing portfolio composed of the underlying asset and a cash account. The portfolio is rebalanced daily to optimally offset the net risk exposure at the option maturity, denoted as $T$ days. The time-$t$ underlying asset price is $S_{t}$. The trading strategy is represented by the predictable process $\delta=\{\delta_{t}\}_{t=1}^{T}$, where $\delta_{t}$ is the number of underlying asset shares held during the interval $(t-1,t]$. The time-$t$ discounted gain made by the hedging portfolio is $G_{t}^{\delta}=\sum_{k=1}^{t}\delta_{k}(\beta^{k}S_{k}\mbox{e}^{q\Lambda}-\beta^{k-1}S_{k-1})$ with $\beta=\mbox{e}^{-r\Lambda}$, where $r$ is the annualized continuously compounded risk-free rate, $q$ is the annualized underlying asset dividend yield, and the period length is $\Lambda=\frac{1}{252}$ years. The time-$t$ self-financing portfolio value is
@@
where $\rho$ is the risk measure used by the agent to quantify risk. In this paper we consider the Conditional Value-at-Risk (CVaRα) defined as $\rho(\xi_{T}^{\delta})=\mathbb{E}[\xi_{T}^{\delta}\mid\xi_{T}^{\delta}\geq\mbox{VaR}_{\alpha}(\xi_{T}^{\delta})]$, where $\alpha\in(0,1)$ and $\mbox{VaR}_{\alpha}(\xi_{T}^{\delta})$ is the Value-at-Risk defined as $\mbox{VaR}_{\alpha}(\xi_{T}^{\delta})=\min_{c}\{c:\mathbb{P}(\xi_{T}^{\delta}\leq c)\geq\alpha\}$. The CVaR is a commonly used objective function in the deep hedging literature, see for instance Carbonneau and Godin, (2021), Cao et al., (2023) or Wu and Jaimungal, (2023). In addition, an appealing feature of the CVaR is that it allows to finetune the investor’s attitude towards risk through the confidence level. A high value of $\alpha$ puts more emphasis on risk reduction, whereas a low value of $\alpha$ penalizes losses and rewards gains.
@@ 3.1   Deep hedging
The deep hedging (DH) framework, introduced by Buehler et al., (2019), provides a solution to the hedging problem (3) by leveraging RL techniques. The DH policy $\tilde{\delta}$ is approximated with a neural network $\delta^{DH}_{\theta}$ with parameters $\theta$, which returns a hedging position $\delta_{t+1}$ when provided with time-$t$ input features $X_{t}$. The objective function to be minimized is thus
@@
which ensures that the cash amount borrowed in the portfolio remains below $B>0$ (see François et al., (2024)).
@@ 3.3   Statistical arbitrage
that is, if the trading strategy $\delta$ which requires zero initial investment is deemed strictly less risky than a null investment according to risk measure $\rho$. Such definition is also in line with that of Buehler et al., (2021), who focus on the case of the entropy risk measure.
@@
is a statistical arbitrage strategy according to risk measure $\rho$. We wish to further extend their study and examine if the trading strategy $\delta^{-}$ behaves like a statistical arbitrage in more general incomplete market dynamics, using a GARCH-based market model as a representative candidate for illustration. In other words, we investigate whether the deep hedging approach typically incorporates a speculative arbitrage-like component aimed at exploiting the structure of the risk measure considered.
@@ 4.1   Stochastic market dynamics
where $\mu,\,\gamma\in\mathbb{R}$, $\omega,\,\alpha,\,\beta$ are positive, $\mathds{1}_{A}$ is the dummy variable indicating if event $A$ occurs and $\{\epsilon_{t}\}_{t=1}^{T}$ are independent standard normal random variables. Parameter estimates are obtained through maximum likelihood on a daily time series of the S&P 500 index extending from January 4, 2016, to December 31, 2020. Estimated parameters are $\mu=0.06\%$, $\omega=0.01\%$, $\alpha=0.11$, $\gamma=0.20$ and $\beta=0.78$. Furthermore, in all experiments, the annualized continuously compounded risk-free rate and dividend yield are assumed to be constant with values set to $r=2.66\%$ and $q=1.77\%$, respectively. These values represent the historical averages of the 1-year zero-coupon yield and the annualized S&P 500 dividend yield over the period extending from 1996 to 2020.
@@ 4.2   Comparative analysis of deep hedging and delta hedging strategies
In this section, we study the relationship between delta hedging and deep hedging. More specifically, we investigate whether the difference between both strategies represents a speculative overlay reminiscent of a statistical arbitrage. The comparison is conducted by hedging the at-the-money (ATM) European call option, with $S_{0}=K=100$, and maturity $T=63$ days. The leverage constraint is $B=100$.
@@
For all confidence levels $\alpha$ below $50\%$, the strategy $\delta^{-}$ exhibits both positive average profitability $\mathbb{E}[V_{T}^{\delta^{-}}(0)]$ and a CVaR value $\rho(-V_{T}^{\delta^{-}}(0))$ that is negative. This corresponds to a formal statistical arbitrage strategy. Moreover, for the case $\alpha=50\%$, even if the risk measure is positive, it is nevertheless negligible in comparison to expected profits. The strategy $\delta^{-}$, though not a statistical arbitrage from the definition, exhibits a behavior that is quite similar to a statistical arbitrage. This might indicate that the Buehler et al., (2021) approach, which consists in using a change of measure under which statistical arbitrage strategies are removed, might be insufficient to prevent speculative behavior from the hedging agent; such speculative strategies do not always qualify as formal statistical arbitrages. Conversely, difference strategies $\delta^{-}$ using $\alpha\geq 85\%$ clearly do not qualify as statistical arbitrage; the associated risk measure is high and the average profitability is negative.
@@
  • Results are for a short position on the ATM call option with a maturity of $N=63$ days. They are computed using 100,000 out-of-sample paths. The metric $\varrho$ denotes the (unconditional) Spearman correlation between underlying asset positions of the delta hedging strategy and the deep hedging strategy across all rebalancing days while $R^{2}$ represents the $R^{2}$ statistic obtained after regressing deep hedging positions onto delta hedging positions.
@@ 5   Conclusion
The two main conclusions from this study are therefore that (i) the objective function of the deep hedging problem must be carefully selected to prevent the hedging agent from abandoning its hedging objective and pursuing speculative behavior, and (ii) deep hedging can soundly achieve its hedging objectives when provided with a suitable risk measure. A possibility could be to use risk measures that do not provide any reward for gains, such as the semi-RMSE used in Carbonneau and Godin, (2023). However, this would come with the cost of negatively impacting the profitability of the strategy. More research is therefore required to determine what risk measure could be used in the objective function to produce sound hedging behavior.